OpenStudy (minecraftqueen):

William put $100 in a bank account last year. Now he has 1/12 more money in the account. Which expressions represent how much money William has in his bank account now? Choose all answers that are correct. A.$100-(1/12x$100) B.$100+(1/12x$100) C.$100+(13/12x$100) D.13/12x$100

OpenStudy (inowalst):

You're looking for addition. "Now he has 1/12 more money" You want to multiply 100 times 1/12 to see how much he has now. What do you think?

OpenStudy (minecraftqueen):

b and c

OpenStudy (inowalst):

Okay, good. However for C we have a "13" that isn't part of the problem, so we can cancel that one out. B is correct.

OpenStudy (anonymous):

OOOOOH a 13. The evil number.

OpenStudy (minecraftqueen):

what dose that have to do with the question just asking ?

OpenStudy (inowalst):

@minecraftqueen Just ignore.

OpenStudy (anonymous):

Someone's jealous I see.

OpenStudy (inowalst):

@hartnn @Abhisar

OpenStudy (anonymous):

What did I do?

OpenStudy (inowalst):

You're posting unrelated things.

OpenStudy (anonymous):

And the question was already answered.

hartnn (hartnn):

how about option D too ?

OpenStudy (inowalst):

I was thinking why the 13 was there.

hartnn (hartnn):

1+1/12 = (12+1)/12 = 13/12

OpenStudy (inowalst):

Oh okay.

hartnn (hartnn):

so B and D ^_^

OpenStudy (minecraftqueen):

thx so much

hartnn (hartnn):

welcome ^_^

OpenStudy (igreen):

Lol